OpenStudy (anonymous):

Find all solutions in the interval [0, 2π). tan x + sec x = 1

OpenStudy (anonymous):

x = 0 x = pi/4 No solution x = 5pi/4

OpenStudy (anonymous):

-1+tanx+secx=0 y = tan(x/2), sin(x) = 2y/(y^2+1), cosx = (1-y^2/y^2+1) Making it into a rational equations -2y/(y-1) = 0 y = 0 x/2 = pin \[x=2 \pi n\ for \ n \in Z\]

OpenStudy (anonymous):

I got pi/4 @iambatman

OpenStudy (anonymous):

@dumbcow

OpenStudy (anonymous):

either that or 5pi/4

OpenStudy (dumbcow):

x = 0, 2pi, 4pi etc since domain is (0,2pi) , answer must be 0